### v1

@@ 3.4 System and Hardware Contexts
Systems determine orchestration; hardware sets feasibility. Pure QML maximizes coherence yet faces trainability limits. Rotation-equivariant ansätze curb barren plateaus [69]; QGANs achieve strong fidelity with hardware-aware scheduling and mitigation [70]; and quantum neural states support block-local edits [71]. In QFU, modular circuits and block partitioning enable selective reset or reversal with classical controllers managing proofs and audits [69, 71]. Temporal hybrids like QK-LSTM embed sequences in quantum feature spaces and converge reliably across rounds [50]. Topology-aware orchestration reduces wall-clock unlearning time without sacrificing fidelity [72].
Hardware choices shape the menu of mechanisms. Simulators offer reproducibility and controlled ablations; QRDR provides logarithmic scaling for compression [73]. Superconducting QPUs deliver fast parameterized gates and low-latency resets [74]. Trapped ions sustain deep coherence for high-fidelity inversion and noise-precision calibration [75]. Annealers implement unlearning as Hamiltonian reconfiguration by flipping local fields [76]. Deep variational circuits with curvature-aware optimizers support structured forgetting in vision workloads [77]. These options define a practical design space in which guarantees and mechanisms must be matched to device realities.
@@ 4.1 Influence- and QFI-weighted unlearning
The goal is simple to state and hard to achieve: remove the effect of $D_{r}$ while preserving utility on $D_{s}=D\setminus D_{r}$. Classically, the first-order impact of deleting $x_{i}$ at an optimum $\theta^{\star}$ is well captured by the influence-function step $\Delta\theta_{i}\approx-H_{\theta^{\star}}^{-1}\nabla_{\theta}L(\theta^{\star};x_{i})$ [16]. This view supports certified corrections that bound the gap to a counterfactual retrain [17, 2]. In the quantum setting, PQCs expose exact hardware gradients via the parameter-shift rule [24], so per-sample or per-client gradients are measurable. Replacing $H^{-1}$ with the inverse QFI aligns the step with the local geometry and improves stability on NISQ devices [28]:
@@
for mini-batches $B\subseteq D_{r}$. Sensitivity governs effectiveness. Local QFI spectra and effective dimension shape generalization [78]. Expressivity and gate choices modulate gradient scales and plateau risk [79]. Layerwise growth preserves gradient signal [80], while dissipative VQAs mix reset and stochastic gates to steer toward low-sensitivity regions [81]. Beyond variational learning, reservoirs move adaptation to a linear readout where marked samples can be down-weighted [82]. Training “rewinding” can undo influential segments using Fisher analysis [83]. Graph-partitioned control exposes modular subcircuits for local edits [84]. QCNN encoders stabilize gradients so that forgetting can target the encoder [85]. Hybrid encoders amplify structure in small data and enable edits in embedding space [86]. In federated settings, these steps compose with masked sharing and topology-aware coordination [87].
@@ 4.2 Parameter reinitialization and partial retraining
Reinitialization localizes change while protecting the useful hypothesis. Partition $\theta=(\theta_{\rm keep},\theta_{\rm forget})$, reset the latter from a reference $p_{0}$, and fine-tune a small set $\Theta(\mathcal{S})$ on $D_{s}$ with a stabilizer $R$ that preserves observables or kernels. Where to reset depends on where information lives. Quantum kernels suggest most influence sits in the feature map; changing maps can outperform expensive kernel retraining [88]. Memetic and architecture search reveal sensitive groups [89]. Quantum pointwise convolutions provide natural reset points for cross-channel mixing [90]. Hybrid QNNs often match classical accuracy with far fewer parameters, so quantum-layer resets yield large effect with small retraining [91]. Model families admit tailored edits: QRBMs and generators reset priors/latents while keeping critics [92, 93]; graph and molecular pipelines reset encoders while preserving heads [94, 95]; hybrid quantum–neural wavefunctions reset quantum blocks and use a classical corrector [96].
A geometric view clarifies why this works. With QFI $F(\theta)$, a local surrogate for the counterfactual gap is $\delta(\theta)\approx\|\theta-\theta^{\setminus D_{r}}\|_{F(\theta)}^{2}$. Resetting coordinates with large QFI eigenvalues maximizes forgetting per parameter. Natural-gradient fine-tuning within $\Theta(\mathcal{S})$ then recovers utility without leaving the unlearning manifold. We monitor trace distance or infidelity; for kernels, MMD and alignment provide efficient surrogates [88]. For generators, provenance signals from hardware noise help attest that the post-unlearning model is not the pre-unlearning instance, and privacy audits probe membership and inversion risk [93]. These diagnostics connect directly to the reporting checklist in Sec. 5.1.
@@ 4.3 Empirical and certified forgetting
Evidence comes in two forms. Empirical tests probe membership inference on $D_{r}$, accuracy on held-out $D_{r}$, and retention on $D_{s}$. Large QML studies show mixed trends and strong dependence on feature maps [97, 98]. Noise can aid forgetting through contractivity [99]. Even compact Lorentz-equivariant QGNNs, while accurate, need removal tests under shift [100].
@@
Classical tools use stability, low-rank corrections [17], SISA partitioning [2], or deletion-aware sampling [23]. In QML, measurement design tightens generalization [101]; QFI spectra and effective dimension supply geometry-aware bounds that complement (6). Kernel surrogates yield empirical certificates: if $K^{\prime}$ aligns with $\bar{K}$ while suppressing $D_{r}$-specific structure, loss correlates with the alignment gap [88]. Hybrid QEKLR separates feature maps from convex readouts for clean post-hoc checks [62]. Explainability strengthens audits: Shapley values reveal block importance [63]; drops in targeted blocks support removal claims. Reverse engineering of transpiled circuits remains a risk [102], so cryptographic provenance is valuable. In federated settings, resilience matters. Auction-based selection and ring topologies harden aggregation [103]. Certified statements can hold at the aggregate when client updates meet Lipschitz-like conditions, which aligns with QFL evidence [98, 97]. These certification practices foreshadow the formal and ethical agenda in Sec. 6.
@@ 4.4 QFL architectures and their unlearning levers
Architecture dictates leakage channels and the cost of edits. In the model plane, kernel QFL shares feature maps and aggregates readouts, often preferable to heavy retraining [88]. Near-optimal kernel PCA scales dimensionality reduction [104]. In oncology, lower-expressivity encodings prove more robust on hardware [105]. Variational QFL trains PQCs or hybrids: QAEs yield compact features [106]; QCNN hybrids improve NISQ feature learning [107]; topology-first search finds circuits before parameter refinement [108]; and QAOA depth motivates depth-aware designs [109]. Reservoir and generative QFL reduce on-hardware backprop and exploit noise shaping [110, 111].
Topology and aggregation complete the picture. Clients update $\theta_{i}$ locally and an aggregator forms $\theta^{t+1}$ from $\{\theta_{i}^{t+1}\}$. Rings reduce central memory and match emerging quantum networks [21]; chain-based QFL removes the server [112]; slimmable QFL adapts width and measurement to channels [113]; CC-QFL uses shadow tomography with classical clients [114]; FedQLSTM accelerates temporal tasks [115]. Privacy levers include gradient hiding over quantum channels [11] and PQFL with homomorphic encryption and QDP [13]. Non-sequential hybrids such as QPIE transfer classical knowledge to flatten Fisher spectra [116]; provenance signals aid audits in generative settings [93]. Architectures that concentrate influence into identifiable modules make unlearning cheaper: feature maps and encoders for kernels and PQCs, readouts for reservoirs, and priors for QGAN/QRBM. These levers connect directly to the scope and mechanism choices of Sec. 3 and to privacy accounting in Sec. 4.6.
@@ 4.5 Client- and entanglement-level forgetting
Entanglement-aware levers support this step. Encoder resets reinitialize feature maps and measurements with light server-side tuning. Feature-space mapping links expressivity to encoding and entanglement patterns [117, 118]. MERA-like pruning collapses correlations [119]. Neuromorphic perceptrons thin entanglement and mitigate plateaus [120]. Concrete handles exist across families: drop client Gram blocks and re-embed with lower-expressivity encodings in kernels [105]; localize influence in encoders for targeted resets in PQC hybrids [106, 107]; forget at the readout in reservoirs [110]; remove latent modes in qGANs with adaptive training [111]; and harden outputs with DP before cryptographic use [121]. Audits should report membership risk for $D_{c}$, alignment shifts, entanglement reduction, and observable norms linked to generalization [122]. Qudits and mixed registers alter crosstalk and twirling costs, so client-local subspaces should be chosen with tracing efficiency in mind [123]. These practices foreshadow fairness and governance issues in Sec. 6.
@@ 4.6 Secure aggregation and differential privacy
and quantum post-processing preserves DP. For QFL, options include homomorphic encryption with quantum updates (PQFL) [13], quantum gradient hiding with low communication [11], and ring/chain topologies that avoid central aggregation [21, 112]. Middleware can align circuit cutting with secure rounds [51]. For PQCs, clip parameter-shift gradients and add Gaussian noise with $\sigma\!\propto\!C\sqrt{2\log(1.25/\delta)}/\varepsilon$; measurement design reduces sensitivity and tightens bounds [101]. For kernels, fix maps, clip clients, and privatize Gram aggregation [88]. Mixed-state classifiers support DP over sufficient statistics [46]. Time-series QFL composes DP across long horizons with a moments accountant [49, 50, 53]. Generative flows and qGANs combine clipping with privatized discriminator statistics and remain robust to adaptive injections [124, 111]. Explainable stacks assist audits [125], and security analytics confirm feasibility on real QPUs with modest parameter counts [52].
A practical round is straightforward: clip, securely aggregate (ring masks or HE), add Gaussian noise, update the moments accountant, and schedule influence-based resets (Sec. 4.2). Dynamic aggregation recovers utility at the same privacy budget [22]. Domain pipelines can embed these steps into scientific circuits [54]. We will report these privacy and audit settings alongside accuracy and physics-aware distances in Sec. 5.1.
@@ 4.7 Complexity, feasibility, and failure modes
Typical failure modes are known. Barren plateaus call for layerwise growth, encoder resets, or dissipative gates [80, 81]. Mis-targeted resets benefit from attributions and kernel alignment [88, 63]. Gradient and kernel leakage is mitigated by clipping, secure aggregation, and DP accounting [11, 13]. Utility collapse on $D_{s}$ is checked by trust-region natural gradients and early stopping. Diagnostics should report membership risk on $D_{r}$ vs. $D_{s}$, probe-set distances in (7), kernel MMD/alignment gaps, entanglement monotones across $(c,\neg c)$, and attribution drops in targeted blocks [88, 63]. QFL audits add per-round $\varepsilon$ and secure-aggregation alarms. These diagnostics and costs will be standardized in Sec. 5.1 and motivate the open challenges in Sec. 6.
@@ 5.1 Metrics and Reporting
  • Capacity and sensitivity. Quantum Fisher information (QFI) spectra and effective dimension quantify model sensitivity to perturbations, guiding geometry-aware unlearning. Measurement Frobenius norms correlate with tighter generalization bounds [101]. Feature attributions such as Shapley values further offer auditable explanations of which circuit blocks carry forgettable information [63].
@@ 5.2 Benchmark Datasets
  • Vision and imaging. Compressed neuroimaging datasets such as CompressedMediQ combine CNN or PCA front-ends with quantum SVM or kernel back-ends [57]. Complex imaging datasets drive VQ-DNN and variational architectures optimized via curvature-aware updates [77].
  • Physics and quantum tasks. Quantum SVM-based entanglement detection benchmarks hardware fidelity across backends [59]. Hamiltonian clustering tasks on annealers offer realistic settings for hybrid optimization and unlearning validation [76].
  • Health, proteomics, and text. Minority-cohort proteomics (QProteoML) test fairness under differential forgetting [58].Fake-news classification using PegasosQSVM includes simulation-to-hardware discrepancy analysis [60]. Temporal health pipelines benchmark recurrent quantum models such as QK-LSTM and QLSTM [50, 49].

### v2

@@ 3.4 System and Hardware Contexts
Systems determine orchestration; hardware sets feasibility. Pure QML maximizes coherence yet faces trainability limits. Rotation-equivariant ansätze curb barren plateaus [sein2025rotation]; QGANs achieve strong fidelity with hardware-aware scheduling and mitigation [pajuhanfard2025qgan]; and quantum neural states support block-local edits [zhang2025nqs]. In QFU, modular circuits and block partitioning enable selective reset or reversal with classical controllers managing proofs and audits [sein2025rotation, zhang2025nqs]. Temporal hybrids like QK-LSTM embed sequences in quantum feature spaces and converge reliably across rounds [hsu2025qklstm]. Topology-aware orchestration reduces wall-clock unlearning time without sacrificing fidelity [phalak2025qualiti].
Hardware choices shape the menu of mechanisms. Simulators offer reproducibility and controlled ablations; QRDR provides logarithmic scaling for compression [yang2025qrdr]. Superconducting QPUs deliver fast parameterized gates and low-latency resets [ding2025pqg]. Trapped ions sustain deep coherence for high-fidelity inversion and noise-precision calibration [bu2025qsp]. Annealers implement unlearning as Hamiltonian reconfiguration by flipping local fields [seong2025hamiltonian]. Deep variational circuits with curvature-aware optimizers support structured forgetting in vision workloads [xu2025vqnn]. These options define a practical design space in which guarantees and mechanisms must be matched to device realities.
@@ 4.1 Influence- and QFI-weighted unlearning
The goal is simple to state and hard to achieve: remove the effect of $D_{r}$ while preserving utility on $D_{s}=D\setminus D_{r}$. Classically, the first-order impact of deleting $x_{i}$ at an optimum $\theta^{\star}$ is well captured by the influence-function step $\Delta\theta_{i}\approx-H_{\theta^{\star}}^{-1}\nabla_{\theta}L(\theta^{\star};x_{i})$ [koh2017understanding]. This view supports certified corrections that bound the gap to a counterfactual retrain [guo2020certified, bourtoule2021machine]. In the quantum setting, PQCs expose exact hardware gradients via the parameter-shift rule [schuld2019evaluating], so per-sample or per-client gradients are measurable. Replacing $H^{-1}$ with the inverse QFI aligns the step with the local geometry and improves stability on NISQ devices [stokes2020quantum]:
@@
for mini-batches $B\subseteq D_{r}$. Sensitivity governs effectiveness. Local QFI spectra and effective dimension shape generalization [khanal2025qfi-bounds]. Expressivity and gate choices modulate gradient scales and plateau risk [liu2025expressibility]. Layerwise growth preserves gradient signal [qi2025dlvqnn], while dissipative VQAs mix reset and stochastic gates to steer toward low-sensitivity regions [ilin2025dvqa]. Beyond variational learning, reservoirs move adaptation to a linear readout where marked samples can be down-weighted [beaulieu2025qrc]. Training “rewinding” can undo influential segments using Fisher analysis [demiranda2025rewinding]. Graph-partitioned control exposes modular subcircuits for local edits [liu2026mpgpqoc]. QCNN encoders stabilize gradients so that forgetting can target the encoder [tang2025qcnn-enc]. Hybrid encoders amplify structure in small data and enable edits in embedding space [jin2025proton]. In federated settings, these steps compose with masked sharing and topology-aware coordination [bhatia2024bhi].
@@ 4.2 Parameter reinitialization and partial retraining
Reinitialization localizes change while protecting the useful hypothesis. Partition $\theta=(\theta_{\rm keep},\theta_{\rm forget})$, reset the latter from a reference $p_{0}$, and fine-tune a small set $\Theta(\mathcal{S})$ on $D_{s}$ with a stabilizer $R$ that preserves observables or kernels. Where to reset depends on where information lives. Quantum kernels suggest most influence sits in the feature map; changing maps can outperform expensive kernel retraining [alvarez2025kernel]. Memetic and architecture search reveal sensitive groups [ardila2025memoqcd]. Quantum pointwise convolutions provide natural reset points for cross-channel mixing [ning2025qpointwise]. Hybrid QNNs often match classical accuracy with far fewer parameters, so quantum-layer resets yield large effect with small retraining [bischof2025entity]. Model families admit tailored edits: QRBMs and generators reset priors/latents while keeping critics [sinno2025qrbm, ghosh2025guardians]; graph and molecular pipelines reset encoders while preserving heads [an2025tensorenc, lu2025qegnns]; hybrid quantum–neural wavefunctions reset quantum blocks and use a classical corrector [li2025hybridwf].
A geometric view clarifies why this works. With QFI $F(\theta)$, a local surrogate for the counterfactual gap is $\delta(\theta)\approx\|\theta-\theta^{\setminus D_{r}}\|_{F(\theta)}^{2}$. Resetting coordinates with large QFI eigenvalues maximizes forgetting per parameter. Natural-gradient fine-tuning within $\Theta(\mathcal{S})$ then recovers utility without leaving the unlearning manifold. We monitor trace distance or infidelity; for kernels, MMD and alignment provide efficient surrogates [alvarez2025kernel]. For generators, provenance signals from hardware noise help attest that the post-unlearning model is not the pre-unlearning instance, and privacy audits probe membership and inversion risk [ghosh2025guardians]. These diagnostics connect directly to the reporting checklist in Sec. 5.1.
@@ 4.3 Empirical and certified forgetting
Evidence comes in two forms. Empirical tests probe membership inference on $D_{r}$, accuracy on held-out $D_{r}$, and retention on $D_{s}$. Large QML studies show mixed trends and strong dependence on feature maps [mohammadisavadkoohi2025survey, qi2025iscas]. Noise can aid forgetting through contractivity [nguyen2020robust]. Even compact Lorentz-equivariant QGNNs, while accurate, need removal tests under shift [jahnin2025lorentz].
@@
Classical tools use stability, low-rank corrections [guo2020certified], SISA partitioning [bourtoule2021machine], or deletion-aware sampling [ginart2019making]. In QML, measurement design tightens generalization [li2025measurement]; QFI spectra and effective dimension supply geometry-aware bounds that complement (6). Kernel surrogates yield empirical certificates: if $K^{\prime}$ aligns with $\bar{K}$ while suppressing $D_{r}$-specific structure, loss correlates with the alignment gap [alvarez2025kernel]. Hybrid QEKLR separates feature maps from convex readouts for clean post-hoc checks [misra2025qeklr]. Explainability strengthens audits: Shapley values reveal block importance [heese2025shapley]; drops in targeted blocks support removal claims. Reverse engineering of transpiled circuits remains a risk [ghosh2025re], so cryptographic provenance is valuable. In federated settings, resilience matters. Auction-based selection and ring topologies harden aggregation [lee2025auction]. Certified statements can hold at the aggregate when client updates meet Lipschitz-like conditions, which aligns with QFL evidence [qi2025iscas, mohammadisavadkoohi2025survey]. These certification practices foreshadow the formal and ethical agenda in Sec. 6.
@@ 4.4 QFL architectures and their unlearning levers
Architecture dictates leakage channels and the cost of edits. In the model plane, kernel QFL shares feature maps and aggregates readouts, often preferable to heavy retraining [alvarez2025kernel]. Near-optimal kernel PCA scales dimensionality reduction [wang2025qkpca]. In oncology, lower-expressivity encodings prove more robust on hardware [repetto2025breast]. Variational QFL trains PQCs or hybrids: QAEs yield compact features [asaoka2025qae]; QCNN hybrids improve NISQ feature learning [li2025cqhcnn]; topology-first search finds circuits before parameter refinement [su2025tdqas]; and QAOA depth motivates depth-aware designs [chicano2025qaoa]. Reservoir and generative QFL reduce on-hardware backprop and exploit noise shaping [ahmed2025qrc, tian2025qgan].
Topology and aggregation complete the picture. Clients update $\theta_{i}$ locally and an aggregator forms $\theta^{t+1}$ from $\{\theta_{i}^{t+1}\}$. Rings reduce central memory and match emerging quantum networks [wang2024quantum]; chain-based QFL removes the server [gurung2025ccqfl]; slimmable QFL adapts width and measurement to channels [park2025slimqfl]; CC-QFL uses shadow tomography with classical clients [song2024ccqfl]; FedQLSTM accelerates temporal tasks [chehimi2024fedqlstm]. Privacy levers include gradient hiding over quantum channels [li2024privacy] and PQFL with homomorphic encryption and QDP [mashetty2025pqfl]. Non-sequential hybrids such as QPIE transfer classical knowledge to flatten Fisher spectra [guo2025qpie]; provenance signals aid audits in generative settings [ghosh2025guardians]. Architectures that concentrate influence into identifiable modules make unlearning cheaper: feature maps and encoders for kernels and PQCs, readouts for reservoirs, and priors for QGAN/QRBM. These levers connect directly to the scope and mechanism choices of Sec. 3 and to privacy accounting in Sec. 4.6.
@@ 4.5 Client- and entanglement-level forgetting
Entanglement-aware levers support this step. Encoder resets reinitialize feature maps and measurements with light server-side tuning. Feature-space mapping links expressivity to encoding and entanglement patterns [walia2025featuremap, hangun2025wind]. MERA-like pruning collapses correlations [zhang2025mera]. Neuromorphic perceptrons thin entanglement and mitigate plateaus [bravo2025qperceptron]. Concrete handles exist across families: drop client Gram blocks and re-embed with lower-expressivity encodings in kernels [repetto2025breast]; localize influence in encoders for targeted resets in PQC hybrids [asaoka2025qae, li2025cqhcnn]; forget at the readout in reservoirs [ahmed2025qrc]; remove latent modes in qGANs with adaptive training [tian2025qgan]; and harden outputs with DP before cryptographic use [ahn2025qryptgen]. Audits should report membership risk for $D_{c}$, alignment shifts, entanglement reduction, and observable norms linked to generalization [hagelueken2025excited]. Qudits and mixed registers alter crosstalk and twirling costs, so client-local subspaces should be chosen with tracing efficiency in mind [venturelli2025qudit]. These practices foreshadow fairness and governance issues in Sec. 6.
@@ 4.6 Secure aggregation and differential privacy
and quantum post-processing preserves DP. For QFL, options include homomorphic encryption with quantum updates (PQFL) [mashetty2025pqfl], quantum gradient hiding with low communication [li2024privacy], and ring/chain topologies that avoid central aggregation [wang2024quantum, gurung2025ccqfl]. Middleware can align circuit cutting with secure rounds [mantha2025pilot]. For PQCs, clip parameter-shift gradients and add Gaussian noise with $\sigma\!\propto\!C\sqrt{2\log(1.25/\delta)}/\varepsilon$; measurement design reduces sensitivity and tightens bounds [li2025measurement]. For kernels, fix maps, clip clients, and privatize Gram aggregation [alvarez2025kernel]. Mixed-state classifiers support DP over sufficient statistics [useche2025qgc]. Time-series QFL composes DP across long horizons with a moments accountant [chen2025dqlstm, hsu2025qklstm, frehner2025qae-ts]. Generative flows and qGANs combine clipping with privatized discriminator statistics and remain robust to adaptive injections [laird2025gflownet, tian2025qgan]. Explainable stacks assist audits [don2025qrlxai], and security analytics confirm feasibility on real QPUs with modest parameter counts [sridevi2025hqc].
A practical round is straightforward: clip, securely aggregate (ring masks or HE), add Gaussian noise, update the moments accountant, and schedule influence-based resets (Sec. 4.2). Dynamic aggregation recovers utility at the same privacy budget [qu2025daqfl]. Domain pipelines can embed these steps into scientific circuits [kang2025mtv]. We will report these privacy and audit settings alongside accuracy and physics-aware distances in Sec. 5.1.
@@ 4.7 Complexity, feasibility, and failure modes
Typical failure modes are known. Barren plateaus call for layerwise growth, encoder resets, or dissipative gates [qi2025dlvqnn, ilin2025dvqa]. Mis-targeted resets benefit from attributions and kernel alignment [alvarez2025kernel, heese2025shapley]. Gradient and kernel leakage is mitigated by clipping, secure aggregation, and DP accounting [li2024privacy, mashetty2025pqfl]. Utility collapse on $D_{s}$ is checked by trust-region natural gradients and early stopping. Diagnostics should report membership risk on $D_{r}$ vs. $D_{s}$, probe-set distances in (7), kernel MMD/alignment gaps, entanglement monotones across $(c,\neg c)$, and attribution drops in targeted blocks [alvarez2025kernel, heese2025shapley]. QFL audits add per-round $\varepsilon$ and secure-aggregation alarms. These diagnostics and costs will be standardized in Sec. 5.1 and motivate the open challenges in Sec. 6.
@@ 5.1 Metrics and Reporting
  • Capacity and sensitivity. Quantum Fisher information (QFI) spectra and effective dimension quantify model sensitivity to perturbations, guiding geometry-aware unlearning. Measurement Frobenius norms correlate with tighter generalization bounds [li2025measurement]. Feature attributions such as Shapley values further offer auditable explanations of which circuit blocks carry forgettable information [heese2025shapley].
@@ 5.2 Benchmark Datasets
  • Vision and imaging. Compressed neuroimaging datasets such as CompressedMediQ combine CNN or PCA front-ends with quantum SVM or kernel back-ends [chen2025compressedmediq]. Complex imaging datasets drive VQ-DNN and variational architectures optimized via curvature-aware updates [xu2025vqnn].
  • Physics and quantum tasks. Quantum SVM-based entanglement detection benchmarks hardware fidelity across backends [mahdian2025qsvment]. Hamiltonian clustering tasks on annealers offer realistic settings for hybrid optimization and unlearning validation [seong2025hamiltonian].
  • Health, proteomics, and text. Minority-cohort proteomics (QProteoML) test fairness under differential forgetting [priyadharshini2025qproteoml].Fake-news classification using PegasosQSVM includes simulation-to-hardware discrepancy analysis [khalil2025pegasosqsvm]. Temporal health pipelines benchmark recurrent quantum models such as QK-LSTM and QLSTM [hsu2025qklstm, chen2025dqlstm].